# MathMist: A Parallel Multilingual Benchmark Dataset for Mathematical Problem Solving and Reasoning

### v1

@@ Abstract
Mathematical reasoning remains one of the most challenging domains for large language models (LLMs), requiring not only linguistic understanding but also structured logical deduction and numerical precision. While recent LLMs demonstrate strong general-purpose reasoning abilities, their mathematical competence across diverse languages remains underexplored. Existing benchmarks primarily focus on English or a narrow subset of high-resource languages, leaving significant gaps in assessing multilingual and cross-lingual mathematical reasoning. To address this, we introduce MathMist  [^1] MathMist Dataset Hugging Face, a parallel multilingual benchmark for mathematical problem solving and reasoning. MathMist encompasses over 21K aligned question–answer pairs across seven languages, representing a balanced coverage of high-, medium-, and low-resource linguistic settings. The dataset captures linguistic variety, multiple types of problem settings, and solution synthesizing capabilities. We systematically evaluate a diverse suite of models, including open-source small and medium LLMs, proprietary systems, and multilingual-reasoning-focused models, under zero-shot, chain-of-thought (CoT), and code-switched reasoning paradigms. Our results reveal persistent deficiencies in LLMs’ ability to perform consistent and interpretable mathematical reasoning across languages, with pronounced degradation in low-resource settings. All the codes and data are available at GitHub: https://github.com/mahbubhimel/MathMist
@@ 1 Introduction
Dataset Lang. Size MCQ CoT CS-CoT Perturb. Reasoning Parallel
@@
MathMist (Ours) 7 21K
Table 1: Comparison of related datasets. Language codes: bn = Bengali, en = English, ar = Arabic, tr = Turkish, ko = Korean, hi = Hindi. CS-CoT = Code-Switched CoT Prompting, Perturb. = Perturbation.
Mathematical Word Problem (MWP) solving requires not only linguistic understanding but also symbolic reasoning, proof formulation, and generalization across typologically diverse languages (Zhang et al., 2020). Yet, existing corpora are either monolingual or translation-based, focusing mainly on final-answer accuracy rather than intermediate reasoning quality (Ahmed et al., 2025; Paul et al., 2025). They have enriched reasoning evaluation, but are confined to arithmetic tasks. Similarly, multilingual datasets expand coverage but lack parallel structure, synthetic perturbations, and multi-format evaluation (Chen et al., 2024). Consequently, there remains a significant gap in understanding how LLMs reason mathematically across different linguistic settings and how they respond to code-switched prompts or fallacious reasoning.
To address these limitations, we introduce MathMist, a parallel multilingual benchmark designed to evaluate LLMs’ mathematical problem solving and reasoning across a diverse set of languages and settings. MathMist provides a controlled platform to assess how models handle linguistic variation, code-switching, and reasoning perturbations within equivalent mathematical contexts. The dataset encompasses over 21K mathematical artifacts spanning seven languages, Bangla, English, French, Finnish, Turkish, Lithuanian, and Kazakh, representing a balanced coverage of high-, medium-, and low-resource languages from distinct linguistic families. Our key contributions are as follows:
  • We introduce MathMist, comprising approximately 21K aligned question–answer pairs across seven languages. The dataset includes 10.1K parallel problems, 2.2K multiple-choice questions, and 8.6K perturbed solutions, enabling fine-grained evaluation of reasoning behaviors and error sensitivity across linguistic settings.
@@ 2 Related Work
Recent research has begun to make measurable progress toward enhancing mathematical reasoning capabilities in low-resource languages. Ahmed et al. (2025) introduced BenNumEval, a benchmark for numerical reasoning in Bengali with six task families and over 3.2k problems, showing that even advanced prompting methods such as Cross-Lingual Prompting (XLP) and Cross-Lingual Chain-of-Thought (XCoT) fall short of human-level performance. Mondal et al. (2025) released BMWP, containing 8,653 Bengali math word problems for operation prediction using deep models, achieving 92% accuracy. Era et al. (2024) created PatiGonit, which contains 10k problems for equation translation and found that transformer models like mT5 reached 97.3% accuracy. Building on these, Paul et al. (2025) introduced SOMADHAN, containing 8,792 manually annotated math word problems with step-by-step reasoning, showing that few-shot Chain-of-Thought prompting improved accuracy up to 88% with LLaMA-3.3-70B. Despite these efforts, most Bengali datasets are limited to arithmetic problems and final-answer evaluation. Beyond Bengali, Mahgoub et al. (2024) proposed a Synthetic Data Augmentation framework for Arabic mathematical problem solving, showing improvements under data scarcity.
@@
Multilingual studies have explored how cross-lingual training and data balance affect reasoning. Chen et al. (2024) introduced MGSM8KInstruct, a dataset spanning 10 languages, demonstrating that multilingual supervised fine-tuning improves both cross-lingual and monolingual reasoning. Zhang et al. (2024) further showed that self-distillation from high-resource to low-resource languages enhances accuracy while reducing data and translation costs. These findings highlight the importance of balanced multilingual corpora and adaptive fine-tuning. Moreover, LLMs are increasingly used for education and tutoring involving mathematical reasoning. Tonga et al. (2025) simulated multilingual LLM tutoring and found that native-language feedback improves reasoning, especially in low-resource contexts. Mahran and Simbeck (2025) developed a multilingual pipeline for generating and grading math problems, revealing persistent linguistic bias favoring English outputs.
Across prior work, low-resource languages have received limited attention. Also, existing datasets focus mainly on arithmetic problems and lack symbolic and proof-based reasoning tasks. They also emphasize final answers rather than intermediate reasoning. Our work addresses these gaps by introducing a comprehensive multilingual mathematical reasoning benchmark with decision-aware, answer-type–specific evaluation, offering fine-grained assessment of reasoning performance. The comparison of mathematical reasoning datasets across multiple languages is shown in Table 1.
@@ 3.1 Data Sourcing
Primarily MathMist contains a total of 2,890 math word problems, with 1,445 problems available in Bangla and English each. The problems were sourced from the National Curriculum & Textbook Board (NCTB) secondary school mathematics books for the 2018–2019 academic year, which had more exemplary content than recent editions. We excluded chapters requiring diagrammatic reasoning from the Mathematics book (chapters 6, 7, 8, and 15) and the Higher Mathematics book (chapters 3, 4, and 11), focusing only on those answerable without illustrations. All included examples and exercises were authored and verified by mathematics experts appointed by the Government of Bangladesh. We intentionally kept variable names and mathematical expressions wrapped up with LaTeX to facilitate future automated processing.
@@ 3.3 Perturbation Generation
In our perturbation-injection pipeline, each item is denoted as $(Q,S_{\text{true}},A)$ with one of three strategies from $\Sigma=\{\sigma_{1},\sigma_{2},\sigma_{3}\}$. The strategy $\sigma_{1}$ infiltrates {step omission, incorrect rule, faulty causality}, $\sigma_{2}$ insinuates {overgeneralization, logical fallacy}, and $\sigma_{3}$ includes all five fallacy types. For each item, we forge $\tilde{S}=\sigma_{i}(S_{\text{true}})$ while ensuring that the perturbed solution keeps the tone and structure of $S_{\text{true}}$, with errors seamlessly embedded and no meta $\mathcal{M}$ tokens present (i.e., $\forall t\in\tilde{S},;t\notin\mathcal{M}$). Each instance $(Q,A,\tilde{S},\text{strategy}=\sigma_{i})$ must pass automated quality checks by a language model. We also manually verified an arbitrary sample $R$ ($|R|\approx 0.2N$) to confirm the efficacy of the error injections. For the English–Bangla parallel corpus, we ensure version-wise equivalence, given $(S_{\text{en}},S_{\text{bn}})$ and $\sigma_{i}$, we generate $(\tilde{S}_{\text{en}},\tilde{S}_{\text{bn}})$ while preserving error types and their locations, enabling fair cross-lingual evaluation.
@@ 3.4 Multilingual Translation Pipeline
To evaluate LLM families on MWP solving across diverse resource levels while ensuring typological variety, we selected high-resource languages such as French (Indo-European); medium-resource languages like Finnish (Uralic) and Turkish (Turkic), and low-resource languages, including Kazakh (Turkic), Lithuanian (Indo-European). With this selection, we augmented our dataset by embodying systematic mathematical translations. Each item in the dataset is denoted as $(Q,S^{\mathrm{true}},A)$. For the target language set $L$ = {French, Kazakh, Finnish, Lithuanian, Turkish}, language-specific translator agents $\mathcal{A}_{\ell}$ generate candidate translations $(Q_{\ell},S_{\ell},A_{\ell})$ using a zero-shot prompt. Each candidate is evaluated by a verifier LLM $V_{\ell}$ based on criteria $\mathcal{Q}=\{\mathcal{M},\mathcal{T},\mathcal{L},\mathcal{C}\}$, which comprise mathematical fidelity, terminological correctness, clarity, and completeness. A candidate is accepted if they meet all criteria. If any standards fails, the verifier $V_{\ell}$ produces a corrected translation $(\tilde{Q_{\ell}},\tilde{S_{\ell}},\tilde{A_{\ell}})=V_{\ell}(Q_{\ell},S_{\ell},A_{\ell};Q,S_{\mathrm{true}},A)$, which must satisfy all the criteria. We evaluated several translations from each language, utilizing back translation, and found acceptable results with minimal distortions. Although the translations are not human-verified, we trust our rigorous pipeline to maintain isomorphism with the original, creating a strong prototype of mathematical reasoning in diverse resource languages.
@@ 3.5 Corpus Statistics
The multilingual corpus consists of 1,445 math problems for each of these languages: Bangla, English, French, Kazakh, Finnish, Lithuanian, and Turkish, totaling 10,115 problem instances. It comprises 5,901 Numerical problems (58.34%), 2,030 Symbolic problems (20.07%), and 2,184 Proof problems (21.59%). Within the Numerical category, arithmetic and algebra are most common, with 2,142 and 1,386 instances, respectively. Furthermore, most Symbolic problems are algebraic, summing 1,610 math problems. Measurements and algebra are dominant in the Proof category. Table 2 shows the statistics for one representative language, as all languages share the same distribution. Additionally, 2,266 multiple-choice questions (MCQs) were created for Bangla and English, along with 8,670 incorrect solution variants. In total, the corpus consists of 21,051 artifacts: 10,115 problems, 2,266 MCQs, and 8,670 perturbed solutions. MathMist facilitates multilingual modeling of problem comprehension and evaluates LLMs’ solving capability.
4 Experimental Setup
4.1 Models
We evaluated state-of-the-art LLMs ranging from 7B–20B parameters, covering both open-source and proprietary families. The assessment spans high- to low-resource languages, varied question structures, and code-switched chain-of-thought reasoning. Our model lineup includes open-source systems such as DeepSeek R1-7B (Guo et al., 2025), Mathstral-7B (Jiang et al., 2023), Qwen-8B (Yang et al., 2025), and GPT-OSS-20B (Agarwal et al., 2025), alongside the proprietary Gemini 2.5 Flash-Lite model for benchmarking against enterprise-grade performance.
Category Subdomain Count
@@ 4.1 Models
Across 7 Languages $1,445\times 7=10,115$
@@
Grand Total Artifacts 21,051
Table 2: Overall corpus statistics across all seven languages in MathMist.
@@ 4.2 Evaluation Metrics
Here, the judge function, $J$, takes LLM’s answer ($A_{\text{llm}}$) and the actual ground truth ($A_{\text{gt}}$) as input. It employs a valuation function, $v$, to convert them to their true mathematical values, expressions, statements, or references. It then estimates whether the distance between them, calculated using the norm $\left\|\cdot\right\|$, falls within a predefined tolerance, $\epsilon$. The function returns a tuple containing thorough reasoning for the decision ($R$) and a binary score (1 for correct, 0 for incorrect). Any cases that cannot be resolved are conservatively categorized as incorrect. We manually inspected 20% of the data for reliability and found all correct. Performance for tasks involving LLMs is evaluated using the Pass@3 metric, which represents success as achieving a solution in at least one of three independent attempts.

### v2

@@ Abstract
Mathematical reasoning remains one of the most challenging domains for large language models (LLMs), requiring not only linguistic understanding but also structured logical deduction and numerical precision. While recent LLMs demonstrate strong general-purpose reasoning abilities, their mathematical competence across diverse languages remains underexplored. Existing benchmarks primarily focus on English or a narrow subset of high-resource languages, leaving significant gaps in assessing multilingual and cross-lingual mathematical reasoning. To address this, we introduce MathMist, a parallel multilingual benchmark for mathematical problem solving and reasoning. MathMist encompasses 2,890 parallel Bangla-English gold standard artifacts, totaling $\approx$30K aligned question–answer pairs across thirteen languages, representing an extensive coverage of high-, medium-, and low-resource linguistic settings. The dataset captures linguistic variety, multiple types of problem settings, and solution synthesizing capabilities. We systematically evaluate a diverse suite of models, including open-source small and medium LLMs, proprietary systems, and multilingual-reasoning-focused models under zero-shot, chain-of-thought (CoT), perturbated reasoning, and code-switched reasoning paradigms. Our results reveal persistent deficiencies in LLMs’ ability to perform consistent and interpretable mathematical reasoning across languages, with pronounced degradation in low-resource settings. All the codes and data are available at GitHub: https://github.com/mahbubhimel/MathMist
@@ 1 Introduction
[…]
@@
[…]
Table 1: Comparison of mathematical reasoning datasets. Language codes: bn = Bengali, en = English, zh = Chinese, ar = Arabic, tr = Turkish, ko = Korean, hi = Hindi. CS-CoT = Code-Switched Chain-of-Thought, Perturb. = Perturbation.
Mathematical Word Problem (MWP) solving requires not only linguistic understanding but also symbolic reasoning, proof formulation, and generalization across typologically diverse languages (Zhang et al., 2020). Yet, existing corpora are either monolingual or fully translation-based, focusing mainly on final-answer accuracy rather than intermediate reasoning quality (Ahmed et al., 2025; Paul et al., 2025). They have enriched reasoning evaluation, but are confined to arithmetic tasks. Similarly, multilingual datasets expand coverage but lack parallel structure, synthetic perturbations, and multi-format evaluation (Chen et al., 2024). Consequently, there remains a significant gap in understanding how LLMs reason mathematically across different linguistic settings and how they respond in code-switched settings or fallacious reasoning.
To address these limitations, we introduce MathMist, a parallel multilingual benchmark designed to evaluate LLMs’ mathematical problem solving and reasoning across a diverse set of languages and task settings. MathMist provides a controlled platform to assess how models handle linguistic variation, code-switching, and reasoning perturbations within equivalent mathematical contexts. The dataset encompasses $\approx$30K mathematical artifacts spanning thirteen languages—English, Arabic, Bangla, French, Swahili, Persian, Turkish, Hausa, Gujarati, Amharic, Kazakh, Finnish, and Lithuanian. This selection covers a wide range of high-, medium-, and low-resource languages from various linguistic families, collectively reaching around 3.15 billion speakers (see Appendix A.3). Our key contributions are as follows:
  • We introduce MathMist, a comprehensive dataset that encompasses 2,890 parallel Bangla-English gold standard math word problems. It comprises approximately 30K aligned question-answer pairs verified by subject matter experts across thirteen languages. The dataset includes 18.9K parallel problems, 2.2K multiple-choice questions, and 8.6K perturbed solutions, enabling fine-grained evaluation of reasoning behaviors and error sensitivity across various linguistic settings.
@@
Overall, MathMist provides the first large-scale, parallel, and linguistically diverse benchmark for mathematical reasoning across multiple languages. By enabling systematic cross-lingual evaluation, it paves the way for a deeper understanding of multilingual reasoning processes in LLMs and establishes a foundation for developing models that reason more reliably, equitably, and transparently across the world’s languages.
@@ 2 Related Work
Recent research has begun to make measurable progress toward enhancing mathematical reasoning capabilities in low-resource languages. Ahmed et al. (2025) introduced BenNumEval, a benchmark for numerical reasoning in Bengali with six task families and over 3.2k problems, showing that even advanced prompting methods such as Cross-Lingual Prompting (XLP) and Cross-Lingual Chain-of-Thought (XCoT) fall short of human-level performance. Mondal et al. (2025) released BMWP, containing 8,653 Bengali math word problems for operation prediction using deep models, achieving 92% accuracy. Era et al. (2024) created PatiGonit, which contains 10k problems for equation translation and found that transformer models like mT5 (Xue et al., 2021) reached 97.3% accuracy. Building on these, Paul et al. (2025) introduced SOMADHAN, containing 8,792 manually annotated math word problems with step-by-step reasoning, showing that few-shot Chain-of-Thought prompting improved accuracy up to 88% with LLaMA-3.3-70B. Despite these efforts, most Bengali datasets are limited to arithmetic problems and final-answer evaluation. Beyond Bengali, Mahgoub et al. (2024) proposed a Synthetic Data Augmentation framework for Arabic mathematical problem solving, showing improvements under data scarcity.
@@
Multilingual studies have explored how cross-lingual training and data balance affect reasoning. Chen et al. (2024) introduced MGSM8KInstruct, a dataset spanning 10 languages, demonstrating that multilingual supervised fine-tuning improves both cross-lingual and monolingual reasoning. Zhang et al. (2024) further showed that self-distillation from high-resource to low-resource languages enhances accuracy while reducing data and translation costs. More recent benchmarks have moved beyond grade-school arithmetic: Luo et al. (2025) proposed MMATH, a multilingual benchmark for complex mathematical reasoning across 10 languages, highlighting language inconsistency and off-target reasoning issues, while Wang et al. (2025) introduced PolyMath to evaluate multilingual mathematical reasoning robustness across diverse linguistic contexts. In parallel, Wu et al. (2024) introduced ConceptMath, a bilingual (English–Chinese) concept-wise benchmark enabling fine-grained analysis of reasoning performance across hierarchical mathematical concepts. These findings highlight the importance of balanced multilingual corpora and adaptive fine-tuning. Moreover, LLMs are increasingly used for education and tutoring involving mathematical reasoning. Tonga et al. (2025) simulated multilingual LLM tutoring and found that native-language feedback improves reasoning, especially in low-resource contexts. Mahran and Simbeck (2025) developed a multilingual pipeline for generating and grading math problems, revealing persistent linguistic bias favoring English outputs.
Across prior work, low-resource languages have received limited attention. Also, existing datasets focus mainly on arithmetic problems and lack symbolic and proof-based reasoning tasks. They also emphasize final answers rather than intermediate reasoning, even in recent multilingual benchmarks (Wu et al., 2024; Luo et al., 2025; Son et al., 2025a). Our work addresses these gaps by introducing a comprehensive multilingual mathematical reasoning benchmark with decision-aware, answer-type–specific evaluation, offering fine-grained assessment of reasoning performance. The comparison of mathematical reasoning datasets across multiple languages is shown in Table 1.
@@ 3.1 Data Sourcing
Primarily, MathMist contains a total of 2,890 gold standard Bangla-English parallel math word problems, with 1,445 problems available in both Bangla and English. The problems were sourced from the National Curriculum & Textbook Board (NCTB) secondary school mathematics books for the 2018–2019 academic year  [^1] https://nctb.gov.bd/ [^2] https://www.panjeree.com/, which had more exemplary content than recent editions. We excluded chapters requiring diagrammatic reasoning from the Mathematics book (chapters 6, 7, 8, and 15) and the Higher Mathematics book (chapters 3, 4, and 11), focusing only on those answerable without illustrations. All included examples and exercises were authored and verified by mathematics experts appointed by the People’s Republic of Bangladesh. We intentionally kept variable names and mathematical expressions wrapped up with LaTeX to facilitate future automated processing.
@@ 3.3 Perturbation Generation
In our perturbation-injection pipeline, each item is denoted as $(Q,S_{\text{true}},A)$ with one of three strategies from $\Sigma=\{\sigma_{1},\sigma_{2},\sigma_{3}\}$, where $S_{\text{true}}$ represents the original correct solution. The strategy $\sigma_{1}$ infiltrates {step omission, incorrect rule, faulty causality}, $\sigma_{2}$ insinuates {overgeneralization, logical fallacy}, and $\sigma_{3}$ includes all five fallacy types. For each item, we forge $\tilde{S}=\sigma_{i}(S_{\text{true}})$ while ensuring that the perturbed solution keeps the tone and structure of $S_{\text{true}}$, with errors seamlessly embedded. Each instance $(Q,A,\tilde{S},\text{strategy}=\sigma_{i})$ must pass automated quality checks by a language model. For the English–Bangla parallel corpus, we ensure version-wise equivalence, given $(S_{\text{en}},S_{\text{bn}})$ and $\sigma_{i}$, we generate $(\tilde{S}_{\text{en}},\tilde{S}_{\text{bn}})$ while preserving error types and their locations. Furthermore, we verified the quality of error injections with subject matter experts (SMEs) to enable fair cross-lingual evaluation.
@@ 3.4 Multilingual Translation Pipeline
To evaluate LLM families on MWP solving across diverse resource levels while ensuring typological variety, we selected high-resource languages such as English, French, and Arabic (Indo-European, Semitic); medium-resource languages like Turkish (Turkic), Persian (Indo-Iranian), Swahili (Bantu), Gujarati (Indo-Aryan), Finnish (Uralic), Lithuanian (Indo-European), and low-resource languages, including Bangla (Indo-Aryan), Hausa (Afroasiatic), Amharic (Semitic), and Kazakh (Turkic). With this selection, we augmented our dataset by embodying systematic mathematical translations. Each item in the dataset is denoted as $(Q,S_{\mathrm{true}},A)$. For the target language set $L$ = {Arabic, French, Swahili, Persian, Turkish, Hausa, Gujarati, Amharic, Kazakh, Finnish, Lithuanian}, language-specific translator agents $\mathcal{A}_{\ell}$ generate candidate translations $(Q_{\ell},S_{\ell},A_{\ell})$ using a zero-shot prompt. Each candidate is evaluated by a verifier LLM $V_{\ell}$ based on criteria $\mathcal{Q}=\{\mathcal{M},\mathcal{T},\mathcal{L},\mathcal{C}\}$, which comprise mathematical fidelity, terminological correctness, clarity, and completeness. A candidate is accepted if they meet all criteria. If any standards fails, the verifier $V_{\ell}$ produces a corrected translation $(\tilde{Q_{\ell}},\tilde{S_{\ell}},\tilde{A_{\ell}})=V_{\ell}(Q_{\ell},S_{\ell},A_{\ell};Q,S_{\mathrm{true}},A)$, which must satisfy all the criteria. To further assess the translation quality, we engaged subject-matter experts (SMEs) for back-translations using Google Translate (Google LLC, 2025) and DeepL (DeepL SE, 2025). If the translations did not match the original English, we provided the candidate translations with explicit error points to Gemini for further refinement. We achieved accurate translations from Gemini on the first attempt due to its strong language coverage. The LLM-SME pipeline maintains isomorphism with the original, creating a robust benchmark MWP dataset for mathematical reasoning across diverse resource languages.
@@ 3.5 Corpus Statistics
The multilingual corpus consists of 1,445 math problems for each of these languages: Bangla, English, French, Kazakh, Finnish, Lithuanian, Turkish, Persian, Arabic, Swahili, Hausa, Gujarati, and Amharic, totaling 18,785 problem instances. It comprises 10,959 Numerical problems (58.34%), 3,770 Symbolic problems (20.07%), and 4,056 Proof problems (21.59%). Within the Numerical category, arithmetic and algebra are most common, with 2,142 and 1,386 instances per language, respectively. Furthermore, most Symbolic problems are algebraic, summing 1,610 math problems per language. Measurements and algebra are dominant in the Proof category. Table 3.5 shows the statistics for one representative language, as all languages share the same distribution. Additionally, 2,266 multiple-choice questions (MCQs) were created for Bangla and English, along with 8,670 incorrect solution variants. In total, the corpus consists of 29,721 artifacts: 18,785 problems, 2,266 MCQs, and 8,670 perturbed solutions. MathMist facilitates multilingual modeling of problem comprehension and evaluates LLMs’ solving capability.
[…]
@@
Across 13 Languages $1,445\times 13=18,785$
@@
[…]
Table 2: Overall corpus statistics across all thirteen languages in MathMist.
4 Experimental Setup
4.1 Models
We evaluated state-of-the-art LLMs ranging from 7B–20B parameters, covering both open-source and proprietary families. The assessment spans high- to low-resource languages, varied question structures, and code-switched chain-of-thought reasoning. Our model lineup includes open-source systems such as DeepSeek R1-7B (Guo et al., 2025), Mathstral-7B (Jiang et al., 2023), Qwen-8B (Yang et al., 2025), and GPT-OSS-20B (Agarwal et al., 2025), alongside the proprietary Gemini 2.5 Flash-Lite model for benchmarking against enterprise-grade performance. To evaluate these models under a standardized experimental setup, we use a decoding temperature of $T=1.0$ to maintain the natural probability distribution and assess calibration. The maximum generation length is set to 32K tokens for long-context reasoning. We also enable "thinking mode" in capable models to capture intermediate reasoning steps.
@@ 4.2 Evaluation Metrics
[…]
Table 3: Scoring rubric and weight distribution for Subject Matter Expert (SME) validation of the LLM-as-a-judge framework on mathematical equivalence tasks.
@@
Here, the judge function, $J$, takes LLM’s answer ($A_{\text{llm}}$) and the actual ground truth ($A_{\text{gt}}$) as input. It employs a valuation function, $v$, to convert them to their true mathematical values, expressions, statements, or references. It then estimates whether the distance between them, calculated using the norm $\left\|\cdot\right\|$, falls within a predefined tolerance, $\epsilon$. The function returns a tuple containing thorough reasoning for the decision ($R$) and a binary score (1 for correct, 0 for incorrect). Any cases that cannot be resolved are conservatively categorized as incorrect. We validate the reliability of our LLM-as-a-Judge framework through SME evaluation. The framework demonstrates high consistency with human judgment, achieving agreement scores ranging from 91.39% to 94.8%. Furthermore, the model exhibits strong qualitative alignment with domain expert reasoning across multilingual mathematical tasks. Table 3 outlines the scoring rubric used in this evaluation, and Appendix A.4 provides additional details. Performance for tasks involving LLMs is evaluated using the Pass@3 metric, which measures whether a problem is solved in at least one of three independent attempts.